Shift summary for Day shift 9/16/99

S. Elizabeth Turner (eturner@kudzu)
Thu, 16 Sep 1999 15:29:54 -0500

Shift Summary for DAY shift, 17/SEP/99

On Shift: Elizabeth, Ugur

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 7.0

Accounting of lost time:

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: None

========================================================================

Accesses during shift: Open but supervised access

========================================================================
Overview of shift: survey crews, AC work, vacuum work, fruit debugging

========================================================================
Description of Shift:

7:25 Survey crew arrives, Hogan LOTOs DCs. DC4 is off.

7:30 Work begins on chilled water -- it gets nice and toasty
in the hall.

8:10 Crew arrives for NM3 vacuum work.

8:30 SET turns off REG transmitter/receiver cards, turns off
MA shaper/splitter, unplugs 1443s for REG from MF15.

Rick C. documents secondary beam changes -- XXI.198

12:25 Reboot ktev3/atev2.

12:30 TRD as founds done -- crew will return to check chamber #1
and re-align if possible.

13:20 Work begins on fan for cooling system.

14:30 NM3 vacuum work complete -- XXI.200

14:50 SET moves MA out of beam.

. . . lots of other stuff that I didn't keep terribly good track of . . .

========================================================================
Follow up on Notes from Previous Shift

Chilled water work done. ET is still off.
DC/TRD surveys done.
NM3 vacuum work done.
REG survey somewhat confused -- will have to be
completed later.
90 degree monitor work done.
MA out.

========================================================================
Items to Follow up on Next Shift

Earliest estimate for beam is 18:00.